OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a green marble? it is 4/19?

OpenStudy (amistre64):

looks right to me; 4 out of 19 for the win

OpenStudy (anonymous):

4/(4+2+5+1+7) =4/19 =21.1% Answer--> 4 out of 19 or 21.1%